Your firm is considering a project that costs $85,105. This opportunity will provide cash flows of $22,311, $10,601, $24,904, and $22,165 over the next four years. Your firm has a required rate of return of 14%. What is the project's profitability index?
Answer #1

profitability index = Present value of cash inflows / initial investment

Present value of cash inflows = 22,311 / (1 + 0.14)1 + 10,601 / (1 + 0.14)2 + 24,904 / (1 + 0.14)3 + 22,165 / (1 + 0.14)4

Present value of cash inflows = 19,571.05263 + 8,157.125269 + 16,809.49064 + 13,123.45935

Present value of cash inflows = 57,661.12789

profitability index = 57,661.12789 / 85,105

profitability index = 0.68